# Multiple Object Detection and Tracking in Panoramic Videos for Cycling Safety Analysis

### v2

@@ Abstract
Cyclists face a disproportionate risk of injury, yet conventional crash records are too limited to reconstruct the circumstances of incidents or to diagnose risk at the finer spatial and temporal detail needed for targeted interventions. Recently, naturalistic studies have gained traction as a way to capture the complex behavioural and infrastructural factors that contribute to crashes. These approaches typically involve the collection and analysis of video data, offering a more detailed and context-rich view of cycling safety. A video promising format is panoramic video, which can record 360° views around a rider. However, its use is limited by severe distortions, large numbers of small objects and boundary continuity, which cannot be handled directly by existing computer vision models. This study addresses these challenges by proposing a novel three-step framework: (1) enhancing object detection accuracy on panoramic imagery by segmenting and projecting the original 360° images into four perspective sub-images, thus reducing distortion; (2) modifying multi-object tracking models to incorporate boundary continuity and object category information for improved tracking consistency; and (3) validating the proposed approach through a real-world application focused on detecting overtaking manoeuvres by vehicles around cyclists. The methodology is evaluated using panoramic videos recorded by cyclists on London’s roadways under diverse conditions. Experimental results demonstrate notable improvements over baseline methods, achieving higher average precision across varying image resolutions. Moreover, the enhanced tracking approach yields a 3.0% increase in multi-object tracking accuracy and a 4.6% improvement in identification F-score. The overtaking detection task achieves a high F-score of 0.81, illustrating the practical effectiveness of the proposed method in real-world cycling safety scenarios. The code is available on GitHub at https://github.com/SpaceTimeLab/360_object_tracking to ensure reproducibility.
@@
journal: Journal name
[…]
@@ 1 Introduction
Cycling, as an active travel mode, has zero carbon emissions (Massink et al., 2011) and can support reductions in traffic congestion (Brunsing, 1997) while improving the health of urban residents (Wanner et al., 2012). Therefore, many local governments and transport authorities are implementing policies to encourage cycling as an alternative to car travel. However, cycling is often perceived as a dangerous activity, particularly in cities that lack high-quality cycling infrastructure. In London, cycle lanes account for only 6.4% of the total road length (excluding highways) (Tait et al., 2022), meaning that cyclists often have to share lanes with motor vehicles. According to Transport for London (2020), despite the mode share of cyclists being only 3.4%, there were 868 killed or seriously injured (KSI) incidents involving cyclists, accounting for 28% of all such casualties in the city in 2020. This disproportionately high risk of incidents can reduce people’s willingness to ride (Delmelle and Delmelle, 2012), which hinders the uptake of cycling. Thus, it is necessary to take measures to improve urban cycling safety, which may include improving road infrastructure (Pucher and Buehler, 2016), implementing policies such as lower speed limits, or running public awareness campaigns like Operation Close Pass [^1] https://www.cyclinguk.org/article/police-close-pass-day-action-tackle-dangerous-overtaking. However, to gain support, these strategies and improvements must be accompanied by evidence of their effectiveness. One aspect of this is the impact on cycling safety.
@@
Although crash data are available through hospital reports and official statistical releases, such as the UK’s STATS19, the low mode share of cycling makes it difficult to draw statistically significant conclusions at a spatially and temporally granular level. Near-miss events, in contrast, are more frequent and have provided a rich source of data for researchers in the form of naturalistic studies (Ibrahim et al., 2021b). These studies have used a range of sensors to collect data while people carry out routine activities on bikes, with the rider indicating when they experience a near-miss. Alternatively, sensors can collect objective data on a near-miss type, such as a close pass (Beck et al., 2019). In recent years, with the rapid development of street view imagery, action/dash cameras, virtual reality (VR) and related technologies, panoramic photography is becoming widespread and mature (Im et al., 2016). Unlike traditional cameras, which can only capture content within a limited field of view, a panoramic camera can capture 180-degree views of the front and rear of the device at the same time with two fisheye lenses, and then concatenate them into one high-resolution 360° image (Hua and Ahuja, 2001). These consumer-grade cameras are suitable to be mounted on helmets or handlebars to record videos of the riders’ surroundings in all directions. Importantly, the 360 view captures the relative movement of the rider and the other people and objects recorded in the scene. For example, in an overtaking manoeuvre, a vehicle will first appear in the rear-view of the camera, before passing into the side view (right or left depending on locality) and the front view. Computer vision based object detection and tracking technologies have the potential to track these movements and categorise them into actions based on interpretations of the relative movements.
@@
The emergence of deep learning has had a transformative effect on a range of machine learning tasks, especially computer vision (LeCun et al., 2015). A range of algorithms have been developed for implementing various visual tasks, such as object detection (Zhao et al., 2019) and multiple object tracking (MOT) (Luo et al., 2021). However, since most of these models and the attached pre-trained weights were designed and trained for traditional images and videos, if applied directly to panoramic videos in equirectangular projection, their performance will be negatively affected (Yang et al., 2018). In addition, some models’ evaluation metrics do not apply to panoramic images. These problems exist because of the following differences between panoramic images and standard images:
  • 1.
@@
  • 3.
    Panoramic images in equirectangular projection have boundary continuity in which the contents of the left and right sides of the image are continuous (Liu et al., 2018).
@@ 2 Related work
Urban transport networks are complex environments in which many agents (pedestrians, cyclists and other road users) interact with each other while using different types of infrastructure. These actions and interactions within the environment contribute to crash risk. Traditionally, crash risk is assessed using historical data on crash records, such as the UK’s STATS19, combined with measures of exposure such as traffic flows, which are combined in a multiple regression framework to identify risk factors (Ambros et al., 2018; Gao et al., 2025). Existing cycling safety studies have heavily relied on crash and conflict data from open government resources (Hels and Orozova-Bekkevold, 2007; Daniels et al., 2008; Zhang et al., 2023), questionnaires surveys (Wang et al., 2020a; Useche et al., 2022), and crowdsourced data reported by volunteer cyclists (Poulos et al., 2012; Fischer et al., 2020). However, these data sources are often subjective and incomplete, making understanding the underlying causes of hazards challenging (Gao et al., 2025). Furthermore, due to the low mode share of cycling in many countries, records on bicycle crashes may be sparse, and exposure data may be limited to particular locations and times or not collected at all.
@@
To complement traditional methods, researchers have turned to naturalistic studies to observe cyclists as they carry out their routine activities (Ibrahim et al., 2021b). Typically, this involves the rider using a suite of sensors that passively collect data such as video, inertial measurement unit (IMU) and location data. Early work relied on customised equipment for this purpose (Johnson et al., 2010), which limited its applicability beyond the scope of the study. Subsequent studies produced sensors for particular purposes, such as detecting close passes, enabling the collection of high-quality, objective data (Beck et al., 2019). Ibrahim et al. (2021b) provide a review of the recent work. While customised equipment allows a high level of control over experimental design, the use of consumer-grade equipment provides an opportunity for larger scale data collection. For example, many road users routinely use dashcams or action cameras to collect data on their journeys to protect themselves if they are involved in a crash. Recognising this, recent work, particularly in autonomous vehicles, has leveraged dashcam video data for crash detection. Rocky et al. (2024) provides a review of the current state of the art. Similar methods have also been applied in cycling safety analysis. For example, using data sourced from YouTube, Ibrahim et al. (2021a) used a deep learning approach that harnessed the optical flow of video data to detect near misses from video streams. Harnessing this data as a form of crowdsourcing could provide a way to monitor safety from the cyclist’s perspective continuously, and studies such as Rick and Bustad (2021) have used popular devices such as GoPro cameras to this end.
@@
The key enabling computer vision technologies for understanding the risk environment from video streams are object detection and multiple object tracking (MOT). Object detection identifies and localises objects, such as vehicles and people, within a video frame. Techniques such as You Only Look Once (YOLO) and Faster Residual Convolutional Neural Network (R-CNN) have been employed that achieve high accuracy in detecting different road users, such as CyDet (Masalov et al., 2018) and MoDeCla (Ojala et al., 2022). MOT is a task that associates multiple objects in different frames of a video into several trajectories (Jiménez-Bravo et al., 2022). MOT technologies have been widely used in the field of transport, particularly in autonomous vehicles. Various modifications have been made to improve car tracking in poor light conditions (Taha et al., 2015), continuously track road users that are sometimes occluded in the videos (Jodoin et al., 2014), and to track objects using the views from multiple cameras (Jiménez-Bravo et al., 2022).
@@
Existing object detection and MOT algorithms are trained on video from cameras with a limited field of view (FOV). This suits the data that has been collected in naturalistic cycling studies to date, which contain only front or rear-facing views. However, risks that cannot be captured in a single view may be present. For example, rear-end crashes have proven to result in a high fatality rate, and tailgating is a risk factor (Australian Transport Safety Bureau, 2006). For a comprehensive view of the risk environment, recording video both in front of and behind the rider is necessary. This can be accomplished using separate cameras, but leads to problems of synchronisation and matching of objects between the two video streams. Furthermore, the side views of the rider are not captured, which are important in the context of certain events such as close passes and car-dooring. Alternatively, panoramic cameras capture 360° views around the user, allowing seamless tracking of interactions between the rider, the environment and other road users.
@@
Existing CV models struggle with panoramic videos due to a lack of panoramic training data, distorted and small objects in equirectangular projections, and issues with boundary continuity. Various studies have attempted to solve these issues. Some have created panoramic datasets to train existing models (Deng et al., 2017; Zhang et al., 2017), while others have applied techniques to improve model performance without retraining (Yang et al., 2018). Still, these do not fully address boundary continuity or the challenge of splitting long objects across subwindows. For MOT, the quality of object detection significantly affects tracking performance. StrongSORT is a commonly used tracker that has been adapted for panoramic data. However, it faces challenges such as object ID switching caused by boundary continuity and inaccuracies in multi-category tracking (MCT). (Liu et al., 2018) proposed extending input images to mitigate boundary continuity issues, yet further improvements are needed to handle objects appearing multiple times across frame boundaries and to enhance MCT.
@@ 3.1 Data collection
To mitigate bias that may be introduced by using video from a single person in a limited range of conditions, validation for overtaking detection is carried out on a set of video clips filmed by 49 different cyclists who were participants in UCL’s 100 Cyclists Project (Christie et al., ). This dataset provides a broader range of riding conditions and rider characteristics on which to test the proposed method: 50 of the videos contain a single clear overtaking event lasting several seconds (‘positive’ examples), and 50 do not (‘negative’). To account for diverse cycling environments, selected videos were captured during daytime (78), night-time (17), and at dawn or dusk (5). 24 videos featured glare from the sun and artificial lighting sources, such as lampposts and headlights. While only one video was recorded in the rain, a further 6 were recorded after recent rain and are slightly obfuscated by raindrops on the camera lens.
@@ 3.2 Data Pre-processing & Annotation
Videos were trimmed and formatted to ‘Horizon Levelling’ by using gyroscope data to level the videos on the horizontal axis. After that, the clips were exported as 5.6k (5,368 × 2,688) equirectangular videos.
To generate labels to evaluate object detection and MOT, the videos were annotated using the Computer Vision Annotation Tool (CVAT) [^2] https://github.com/cvat-ai/cvat. The following 7 categories of objects were chosen for labelling, which are both relevant to cycling safety and contained in the Microsoft COCO dataset (Lin et al., 2014): person, bicycle, car, motorbike, bus, truck and traffic light. In order to output the annotations in both COCO and MOT formats, ‘tracks’ were created to associate the same objects in different image frames.
Following the standards of MOT format, CVAT only supports labelling each object once in each frame. Hence, objects are divided into two parts due to the boundary continuity could be labelled as two separate objects.
To address this, after the annotations were output in both COCO and MOT formats, a method was implemented to merge these split objects in the MOT annotation files without violating the rules of MOT format (Milan et al., 2016), whose workflow is shown in Figure 1. After the ID modifications, 3 videos were split into image frames. These images are used to evaluate object detection models with the COCO annotation files. Finally, after pre-processing and annotation, the dataset has 24,454 labelled objects and 224 labelled tracks in total.
Figure 1: Workflow of merging the split objects in MOT annotation files
@@ 3.3 Exploratory Data Analysis
3.3.1 Position Distribution
To understand where the objects of interest are located, a matrix whose size is the same as the videos (5,368 × 2,688) is used to count how many times each pixel is within an annotated bounding box. Then, the matrix is plotted as a heat map, as Figure 2 shows. As the top and bottom areas of an equirectangular image represent the sky and the ground, respectively, almost all the objects in the dataset are distributed from -50° to 70° of the y axes.
@@ 4.1 Improving Object Detection on Panoramic Images
Figure 3: Workflow of the proposed method for improving object detection on panoramic images
(b) Process of reprojecting an object which is across the boundaries
Figure 4: Projection and aggregation of the detection results of sub-images
Figure 5: Workflow of improved object tracking model using StrongSORT as backbone for better tracking performance on panoramic cycling videos
@@ 4.1.1 Projection Transformation from Equirectangular to Perspective
To present a whole sphere as a 2D image, equirectangular projection maps longitude lines and latitude circles to equally spaced vertical and horizontal lines, respectively (Snyder, 1997). In contrast, perspective projection only maps part of a sphere (within the FOV) from the centre to a tangent plane (Yang et al., 2018). Thus, the transformation from an equirectangular to a perspective is a projection from a sphere to a plane.
According to Lee and Han (2021), assuming that the perspective projection is along the z-axis, the corresponding point P’ $(X,Y,Z)$ on the sphere of the point P $(u,v)$ on the tangent plane can be calculated as:
@@
If represented in a geographic coordinate system, P’ can be also represented as longitude $\phi$ and latitude $\theta$ (Lee and Han, 2021):
@@
If the projection is not towards the z-axis, when calculating the coordinates of point P’ $(X^{\prime},Y^{\prime},Z^{\prime})$, a rotation equation $R$ should be applied as follows (Lee and Han, 2021):
@@
For each output pixel in the perspective projected image, its corresponding location in the original ERP format is computed by Equations (1) to (7). Then the value of the pixel is computed by cubic interpolation where a 4 by 4 block of its neighbouring pixels is involved. To identify parameters for perspective projection transformation, this research referred to the position distribution analysed in subsection 3.3.1: since almost all the objects in the videos are distributed from -50° to 70° vertically, FOV and $\phi_{c}$ were set to 120° and -10°, so that only the objects within this interval would be projected and detected in the improved object detection model.
Although three sub-images with a 120° FOV are enough to cover an equirectangular image horizontally, because there are still some distortions in the perspective images, objects near the edges of the sub-images may not be detected. To solve this problem, the original image is projected into 4 sub-windows along the lines of 0°, 90°, -180° and -90° longitude (i.e., theta) — since there is an overlap of 30° between each pair of adjacent sub-images, objects within 15° from an edge in one sub-image are closer to the centre in another sub-image, which results in less severe distortions.

### v3

@@ Abstract
[Abstract]Cyclists face a disproportionate risk of injury, yet conventional crash records are too sparse to identify risk factors at fine spatial and temporal scales. Recently, naturalistic studies have used video data to capture the complex behavioural and infrastructural risk factors. A promising format is panoramic video, which can record 360 views around a rider. However, its use is limited by distortions, large numbers of small objects and boundary continuity, which cannot be handled using existing computer vision models. This research proposes a novel three-step framework: (1) enhancing object detection accuracy on panoramic imagery by segmenting and projecting the original 360 images into sub-images; (2) modifying multi-object tracking models to incorporate boundary continuity and object category information; and (3) validating through a real-world application of vehicle overtaking detection. The methodology is evaluated using panoramic videos recorded by cyclists on London’s roadways under diverse conditions. Experimental results demonstrate improvements over baselines, achieving higher average precision across varying image resolutions. Moreover, the enhanced tracking approach yields a 10.0% decrease in identification-switches and a 2.7% improvement in identification precision. The overtaking detection task achieves a high F-score of 0.82, illustrating the practical effectiveness of the proposed method in real-world cycling safety scenarios.
@@
computer vision, cycling safety, object detection, object tracking, panoramic videos
articletype: Article Typejournal: Journalvolume: 00footnotetext: The email address of the submitting author: yitai.cheng.21@ucl.ac.uk.footnotetext: *These authors contributed equally to this work.
@@ 1 Introduction
Cycling, as an active travel mode, has zero carbon emissions massink_2011 and can support reductions in traffic congestion brunsing_1997 while improving the health of urban residents wanner2012active. Therefore, many local governments and transport authorities are implementing policies to encourage cycling as an alternative to car travel. However, cycling is often perceived as a dangerous activity, particularly in cities that lack high-quality cycling infrastructure. In London, cycle lanes account for only 6.4% of the total road length (excluding highways) tait2022cycling, meaning that cyclists often have to share lanes with motor vehicles. According to transport_for_london_casualties_2020, despite the mode share of cyclists being only 3.4%, there were 868 killed or seriously injured (KSI) incidents involving cyclists, accounting for 28% of all such casualties in the city in 2020. This disproportionately high risk of incidents can reduce people’s willingness to ride Delmelle2012, which hinders the uptake of cycling. Thus, it is necessary to take measures to improve urban cycling safety, which may include improving road infrastructure pucher2016safer, implementing policies such as lower speed limits, or running public awareness campaigns like Operation Close Passhttps://www.cyclinguk.org/article/police-close-pass-day-action-tackle-dangerous-overtaking. However, to gain support, these strategies and improvements must be accompanied by evidence of their effectiveness. One aspect of this is the impact on cycling safety.
@@
Although crash data are available through hospital reports and official statistical releases, such as the UK’s STATS19, the low mode share of cycling makes it difficult to draw statistically significant conclusions at a spatially and temporally granular level. Near-miss events, in contrast, are more frequent and have provided a rich source of data for researchers in the form of naturalistic studies Ibrahim2021. These studies have used a range of sensors to collect data while people carry out routine activities on bikes, with the rider indicating when they experience a near-miss. Alternatively, sensors can collect objective data on a near-miss type, such as a close pass BECK2019253. In recent years, with the rapid development of street view imagery, action/dash cameras, virtual reality (VR) and related technologies, panoramic photography is becoming widespread and mature Im2016. Unlike traditional cameras, which can only capture content within a limited field of view, a panoramic camera can capture 180-degree views of the front and rear of the device at the same time with two fisheye lenses, and then concatenate them into one high-resolution 360 image hong_hua_2001. These consumer-grade cameras are suitable to be mounted on helmets or handlebars to record videos of the riders’ surroundings in all directions. Importantly, the 360 view captures the relative movement of the rider and the other people and objects recorded in the scene. For example, in an overtaking manoeuvre, a vehicle will first appear in the rear-view of the camera, before passing into the side view (right or left depending on locality) and the front view. Computer vision based object detection and tracking technologies have the potential to track these movements and categorise them into actions based on interpretations of the relative movements.
@@
The emergence of deep learning has had a transformative effect on a range of machine learning tasks, especially computer vision LeCun2015. A range of algorithms have been developed for implementing various visual tasks, such as object detection zhao2019object and multiple object tracking (MOT) Luo2021. However, since most of these models and the attached pre-trained weights were designed and trained for traditional images and videos, if applied directly to panoramic videos in equirectangular projection, their performance will be negatively affected yang2018object. In addition, some models’ evaluation metrics do not apply to panoramic images. These problems exist because of the following differences between panoramic images and standard images:
@@
  • Panoramic images in equirectangular projection have boundary continuity in which the contents of the left and right sides of the image are continuous liu_2018.
@@ 2 Related work
Urban transport networks are complex environments in which many agents (pedestrians, cyclists and other road users) interact with each other while using different types of infrastructure. These actions and interactions within the environment contribute to crash risk. Traditionally, crash risk is assessed using historical data on crash records, such as the UK’s STATS19, combined with measures of exposure such as traffic flows, which are combined in a multiple regression framework to identify risk factors Ambros2018, gao2025reliable. Existing cycling safety studies have heavily relied on crash and conflict data from open government resources Hels2007, Daniels2008, zhang2023analyzing, questionnaires surveys wang2020exploring, useche2022cross, and crowdsourced data reported by volunteer cyclists poulos2012exposure, fischer2020does. However, these data sources are often subjective and incomplete, making understanding the underlying causes of hazards challenging gao2025reliable. Furthermore, due to the low mode share of cycling in many countries, records on bicycle crashes may be sparse, and exposure data may be limited to particular locations and times or not collected at all.
@@
To complement traditional methods, researchers have turned to naturalistic studies to observe cyclists as they carry out their routine activities Ibrahim2021. Typically, this involves the rider using a suite of sensors that passively collect data such as video, inertial measurement unit (IMU) and location data. Early work relied on customised equipment for this purpose johnson_2010, which limited its applicability beyond the scope of the study. Subsequent studies produced sensors for particular purposes, such as detecting close passes, enabling the collection of high-quality, objective data BECK2019253. Ibrahim2021 provide a review of the recent work. While customised equipment allows a high level of control over experimental design, the use of consumer-grade equipment provides an opportunity for larger scale data collection. For example, many road users routinely use dashcams or action cameras to collect data on their journeys to protect themselves if they are involved in a crash. Recognising this, recent work, particularly in autonomous vehicles, has leveraged dashcam video data for crash detection. rocky_review_2024 provides a review of the current state of the art. Similar methods have also been applied in cycling safety analysis. For example, using data sourced from YouTube, ibrahim_cyclingnet_2021 used a deep learning approach that harnessed the optical flow of video data to detect near misses from video streams. Harnessing this data as a form of crowdsourcing could provide a way to monitor safety from the cyclist’s perspective continuously, and studies such as rick_cycling_2021 have used popular devices such as GoPro cameras to this end.
@@
The key enabling computer vision technologies for understanding the risk environment from video streams are object detection and multiple object tracking (MOT). Object detection identifies and localises objects, such as vehicles and people, within a video frame. Techniques such as You Only Look Once (YOLO) and Faster Residual Convolutional Neural Network (R-CNN) have been employed that achieve high accuracy in detecting different road users, such as CyDet masalov_2018 and MoDeCla ojala2022motion. MOT is a task that associates multiple objects in different frames of a video into several trajectories JimnezBravo2022. MOT technologies have been widely used in the field of transport, particularly in autonomous vehicles. Various modifications have been made to improve car tracking in poor light conditions Taha2015, continuously track road users that are sometimes occluded in the videos jodoin_2014, and to track objects using the views from multiple cameras JimnezBravo2022.
@@
Existing object detection and MOT algorithms are trained on video from cameras with a limited field of view (FOV). This suits the data that has been collected in naturalistic cycling studies to date, which contain only front or rear-facing views. However, risks that cannot be captured in a single view may be present. For example, rear-end crashes have proven to result in a high fatality rate, and tailgating is a risk factor australian_transport_safety_bureau_deaths_2006. For a comprehensive view of the risk environment, recording video both in front of and behind the rider is necessary. This can be accomplished using separate cameras, but leads to problems of synchronisation and matching of objects between the two video streams. Furthermore, the side views of the rider are not captured, which are important in the context of certain events such as close passes and car-dooring. Alternatively, panoramic cameras capture 360 views around the user, allowing seamless tracking of interactions between the rider, the environment and other road users.
@@
Existing CV models struggle with panoramic videos due to a lack of panoramic training data, distorted and small objects in equirectangular projections, and issues with boundary continuity. Various studies have attempted to solve these issues. Some have created panoramic datasets to train existing models deng_2017, zhang2017real, while others have applied techniques to improve model performance without retraining yang2018object. Still, these do not fully address boundary continuity or the challenge of splitting long objects across subwindows. For MOT, the quality of object detection significantly affects tracking performance. StrongSORT is a commonly used tracker that has been adapted for panoramic data. However, it faces challenges such as object ID switching caused by boundary continuity and inaccuracies in multi-category tracking (MCT). liu_2018 proposed extending input images to mitigate boundary continuity issues, yet further improvements are needed to handle objects appearing multiple times across frame boundaries and to enhance MCT.
@@ 3.1 Data collection
To mitigate bias that may be introduced by using video from a single person in a limited range of conditions, validation for overtaking detection is carried out on a set of video clips filmed by 49 different cyclists who were participants in UCL’s 100 Cyclists Project christiedetecting. This dataset provides a broader range of riding conditions and rider characteristics on which to test the proposed method: 50 of the videos contain a single clear overtaking event lasting several seconds (‘positive’ examples), and 50 do not (‘negative’). To account for diverse cycling environments, selected videos were captured during daytime (78), night-time (17), and at dawn or dusk (5). 24 videos featured glare from the sun and artificial lighting sources, such as lampposts and headlights. While only one video was recorded in the rain, a further 6 were recorded after recent rain and are slightly obfuscated by raindrops on the camera lens.
@@ 3.2 Data Pre-processing & Annotation
Videos were trimmed and formatted to ‘Horizon Levelling’ by using gyroscope data to level the videos on the horizontal axis. After that, the clips were exported as 5.6k (5,368 $\times$ 2,688) equirectangular videos.
To generate labels to evaluate object detection and MOT, the 3 video clips collected by the research team were annotated using the Computer Vision Annotation Tool (CVAT)https://github.com/cvat-ai/cvat. The following 7 categories of objects were chosen for labelling, which are both relevant to cycling safety and contained in the Microsoft COCO dataset lin_2014: person, bicycle, car, motorbike, bus, truck and traffic light. In order to output the annotations in both COCO and MOT formats, ‘tracks’ were created to associate the same objects in different image frames.
Following the standards of MOT format, CVAT only supports labelling each object once in each frame. Hence, objects are divided into two parts due to the boundary continuity could be labelled as two separate objects. To address this, after the annotations were output in both COCO and MOT formats, a method was implemented to merge these split objects in the MOT annotation files without violating the rules of MOT format milan2016mot16benchmarkmultiobjecttracking. After the ID modifications, videos were split into image frames. These images are used to evaluate object detection models with the COCO annotation files. Finally, after pre-processing and annotation, the dataset has 24,454 labelled objects and 224 labelled tracks in total.
@@ 3.3 Exploratory Data Analysis
To understand where the objects of interest are located, a matrix whose size is the same as the videos (5,368 $\times$ 2,688) is used to count how many times each pixel is within an annotated bounding box. Then, the matrix is plotted as a heat map, as Figure 1 shows. As the top and bottom areas of an equirectangular image represent the sky and the ground, respectively, almost all the objects in the dataset are distributed from -50 to 70 of the y axes.
@@ 4.1 Improving Object Detection on Panoramic Images
Figure 2: Workflow of the proposed method for improving object detection on panoramic images.
@@ 4.1.1 Projection Transformation from Equirectangular to Perspective
To present a whole sphere as a 2D image, equirectangular projection maps longitude lines and latitude circles to equally spaced vertical and horizontal lines, respectively snyder1997flattening. In contrast, perspective projection only maps part of a sphere (within the FOV) from the centre to a tangent plane yang2018object. Thus, the transformation from an equirectangular to a perspective is a projection from a sphere to a plane.
According to Lee2021, assuming that the perspective projection is along the z-axis, the corresponding point P’ $(X,Y,Z)$ on the sphere of the point P $(u,v)$ on the tangent plane can be calculated as:
@@
If represented in a geographic coordinate system, P’ can be also represented as longitude $\phi$ and latitude $\theta$ Lee2021:
@@
If the projection is not towards the z-axis, when calculating the coordinates of point P’ $(X^{\prime},Y^{\prime},Z^{\prime})$, a rotation equation $R$ should be applied as follows Lee2021:
@@
For each output pixel in the perspective projected image, its corresponding location in the original ERP format is computed by Equations (1) to (7). Then the value of the pixel is computed by cubic interpolation where a 4 by 4 block of its neighbouring pixels is involved. To identify parameters for perspective projection transformation, this research referred to the position distribution analysed in subsection 3.3: since almost all the objects in the videos are distributed from -50 to 70 vertically, FOV and $\phi_{c}$ were set to 120 and -10, so that only the objects within this interval would be projected and detected in the improved object detection model.
Although three sub-images with a 120 FOV are enough to cover an equirectangular image horizontally, because there are still some distortions in the perspective images, objects near the edges of the sub-images may not be detected. To solve this problem, the original image is projected into 4 sub-windows along the lines of 0, 90, -180 and -90 longitude (i.e., theta) — since there is an overlap of 30 between each pair of adjacent sub-images, objects within 15 from an edge in one sub-image are closer to the centre in another sub-image, which results in less severe distortions.
(b) Process of reprojecting an object which is across the boundaries
Figure 3: Projection and aggregation of the detection results of sub-images.